OpenStudy (anonymous):

Find the vertical asymptotes, if any, of the graph of the rational function g(x) = x/ x - 5 A) x=0 and x = -5 B) x= 5 C) x = 0 and x = 5 D) no vertical symptote

OpenStudy (callisto):

You have a vertical asymptotes when the graph is undefined at a certain point.. For this question, since it is a fraction and it would be undefined when the denominator is equal to zero. So, put denominator = 0 to find the vertical asymptotes, i.e. x-5 = 0 <- solve this.

OpenStudy (anonymous):

x= 5

OpenStudy (callisto):

That's it, I think.

OpenStudy (anonymous):

I'm confused about B or C that's why i ask

OpenStudy (anonymous):

0 is not included then?

OpenStudy (anonymous):

according to einsteins therory of 12<1678$/9090_+b=b342, you forgot to carry the 9 and the 12 and you have to divide by 70/3

OpenStudy (callisto):

0 is not included. since at x = 0, g(0) = 0/(0-5) = 0, the function is still defined. Got it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

got it THANK YOU

OpenStudy (callisto):

Welcome!!~